Grade 11Physical Chemistry

Calculate the weight of V2O5 produced from 2g of VO & 5.75 g of Fe2O3.

Kshitij Ingle
10 Years agoGrade 11

2 Answers

SAHIL
10 Years ago

2VO + 3Fe2O3 → 6FeO + 2V2O5

molec wt of VO= 67

MW of Fe2O3= 160

MW of V2O5= 182

let weight of V2O5 be x

then we can see 1 mole of V2O5 is formed from 2 moles of VO and 3 moles of Fe2O3

thus x182 = min{22X67,5.753X160}

mohan
8 Years ago
acc. to balance rxn 3 mole fe2o3 form 6 mole feo5.75/160 mole fe2o3 will form 5.169g feototal mass of reactant is 2+5.75=7.75Mass of v2o5=mass of product - mass of feo =mass of reactant-mass of feo =7.750-5.169 =2.581g.